in a rectangle, the ratio of the width to the lengh is 5 : 12 the length is 48 cm. find the width
Answer by Jk22(389)

for twelve units of length, the width is five, if the length is forty-eight centimeter, which is four times twelve, the width would be four times five, hence twenty centimeter.
(w/l=5/12, w=5/12*l=5/12*48=5*4=20 cm)